OpenStudy (anonymous):

Larry's time to travel 357 miles is 3 hours more than Terrell's time to travel 220 miles. Terrell drove 4 miles per hour faster than Larry. How fast did each one travel?

OpenStudy (jdoe0001):

so, if terrell traveled "h" hours, larry traveled "h+3" terrel drove 4 miles faster so if larry drove "r" m/h fast, terrell drove "r+4" m/h faster

OpenStudy (jdoe0001):

so $$ \begin{matrix} & d & t & r \\ \hline larry & 357 & h+3 & r \\ terrell & 220 & h & r+4\\ \end{matrix} $$

OpenStudy (jdoe0001):

so, using the d = rt distance formula larry traveled "t" distance, or t = d/r or \(h+3 = \cfrac{357}{r} \implies h = \cfrac{357}{r}-3\)

OpenStudy (jdoe0001):

terrell's rate is r+4 which \(r+4 =\cfrac{220}{h}\) but \( h = \cfrac{357}{r}-3\) from the 1st equation so $$ r+4 =\cfrac{220}{\cfrac{357}{r}-3} $$

OpenStudy (jdoe0001):

$$ r+4 =\cfrac{220}{\cfrac{357}{r}-3} \implies r+4 = \cfrac{220r}{357-3r}\\ \implies (r+4)(357-3r) = 220r\\ \implies -3r^2-12r+357r+1428-220r = 0\\ \implies \color{blue}{-3r^2+125r+1428=0}\\ \text{from there you'd need to use the quadratic formula}\\ \text{to get "r" or "larry's rate"} $$

OpenStudy (jdoe0001):

@inprogress you'd get 2 values, discard the negative one, since "rate" is positive

OpenStudy (jdoe0001):

@inprogress $$ \text{quadratic formula for it}\\ \cfrac{-125\sqrt{15625+17136}}{-6}\\ \cfrac{-125\pm 181}{-6} $$

OpenStudy (anonymous):

A solution using Mathematica to perform the calculations is attached.